### v1

# CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes

Abstract
We present CD-NGP, which is a fast and scalable representation for 3D reconstruction and novel view synthesis in dynamic scenes. Inspired by continual learning, our method first segments input videos into multiple chunks, followed by training the model chunk by chunk, and finally, fuses features of the first branch and subsequent branches. Experiments on the prevailing DyNeRF dataset demonstrate that our proposed novel representation reaches a great balance between memory consumption, model size, training speed, and rendering quality. Specifically, our method consumes $85\%$ less training memory ($<14$GB) than offline methods and requires significantly lower streaming bandwidth ($<0.4$MB/frame) than other online alternatives. Furthermore, we conduct experiments on a self-collected extra-long video sequences dataset, which demonstrates the superior scalability of our method over previous state-of-the-art approaches.
Keywords:
Continual learning Dynamic novel view synthesis Neural Radiance Field.
1 Introduction
Neural radiance Field (NeRF) [30] has made significant progress in 3D scene reconstruction and novel view synthesis. It takes multi-view RGB images with their camera poses as input and leverages multi-layer perceptrons (MLPs) to reconstruct both the geometry and color of the scene. Many variants have been proposed to improve NeRF’s efficiency [9, 38, 31, 56, 53], reconstruction quality [2, 3, 4], scalability [47, 50], and pose robustness [5, 26, 13, 52, 32, 51]. Recently, 3D Gaussian Splatting (3DGS) [22] proposes an explicit representation with differentiable rendering, which significantly speeds up the reconstruction and enables real-time novel view synthesis. Despite significant progress having been made in this field, the above-mentioned methods assume that the scene is static and may not work properly in dynamic scenes.
To address the dynamic reconstruction problem, amount of efforts have been made under the offline setting, requiring access to the entire dataset in the whole training stage. For example, [36, 28, 34, 33, 15, 29, 44] propose to use MLP-based representation and achieve high-quality results. [40, 17, 7, 16, 45] speed up the optimization by using tensor factorization and voxel representation, respectively. The more recent continual-learning-based methods [27, 48] can train models frame by frame, and streamable representations [43] enable the model to be loaded on the fly at inference. However, the offline methods require huge host memory e.g., taking more than $100GB$ when reconstructing multi-view dynamic videos of 300 frames, and the continual learning-based methods are weak in balancing the convergence speed and model compactness. As a result, these methods are limited in scalability, making them unsuitable for long videos.
In this paper, we propose Continual Dynamic neural graphics primitives (CD-NGP), a novel, fast, and scalable representation for reconstructing dynamic scenes, which enables dynamic reconstruction in long videos. As illustrated in Fig. 1, we evenly split the dynamic videos into multiple chunks based on the timestamp and conduct reconstruction with different model branches, where each branch includes MLP regressors, spatial encoding for $(x,y,z)$, and temporal encoding for $t$. We encode the different chunks by hash tables with different complexities to ensure model scalability and then continually train the model chunk by chunk to alleviate memory occupation. For each individual chunk, we locate the corresponding model branch, sample 3D points $(x,y,z)$ according to frame pixels and cached occupancy grids [31], and represent spatial features by spatial hash grids. The spatial features of the first branch and the current branch are fused by element-wise sum as default or concatenation if applicable. The temporal features are encoded by the hash table for time stamp $t$ in the current branch only. Spatial and temporal features are concatenated along with viewing direction $d$ and input to MLP regressors for rendering and optimization.
Extensive experiments demonstrate the great scalability of our method: compared with prevailing offline baselines, our method consumes $85\%$ less host memory at the training phase, $40\%$ less model size at the inference phase, and delivers comparable results. Compared with online baselines, our method achieves compact model size, high quality, and fast convergence simultaneously. The main contribution of this work is summarized as 3 folds:
Figure 1: The demonstration of our continual learning pipeline for dynamic scenes. Please note the dimensions of representations for different branches remain the same while the complexities of hash tables are flexible and could be configured to ensure scalability. The spatial hash features could be encoded along voxels [31], planes [17] or both [38].
  • We propose a novel method combining spatial and temporal hash encodings for dynamic scenes in short intervals with high scalability.
  • We propose a fast continual learning framework for dynamic scenes, producing comparable results to offline state-of-the-art methods while reducing memory and storage occupation significantly.
  • We provide a dataset of extra-long video sequences ($>3$ minutes) and further demonstrate the scalability of our method on long multi-view videos.
2 Related Work
2.1 Fast novel view synthesis in static scenes
The acceleration of NeRFs falls into two categories: training acceleration and rendering acceleration. Plenoxels [18] uses voxel grids to represent the scene and achieves magnitudes of acceleration over MLP backbones. TensoRF [9] decomposes the 3D scene as the product of tensors and achieves fast novel view synthesis. Plane-based methods [17, 21] project 3D points onto 3 orthogonal 2D planes to ensure sparsity and achieve fast convergence. Instant-NGP [31] leverages an occupancy grid and a multi-resolution hash encoding along with optimized tiny MLPs to enable fast novel view synthesis within a minute. The recent 3D Gaussian method [22] represents the scenes as a union of explicit Gaussian point clouds and achieves high fidelity, fast convergence, and real-time rendering on high-end hardware by eliminating the querying burden of spatial representations. KiloNeRF [37] distills the large MLP in NeRF [30] into tiny MLPs to accelerate rendering by thousands of times. Hedman et al. [20] propose the baking method to accelerate the rendering process. They alpha-compose the cached features along the ray and use MLPs only once per ray to overcome the computation load of the per-point MLP querying in NeRF. BakedSDF [56] stores the view-dependent features into Gaussian spheres along the reconstructed surfaces, and achieves real-time rendering on ordinary laptops. MobileNeRF[10] represents scenes into meshes and enables real-time rendering on mobile devices. Light field methods [42, 8, 19] directly represent the scene as a set of rays to eliminate volume rendering and enable real-time applications on mobile devices.
2.2 Representations of dynamic scenes
There are mainly 3 variants of NeRF for dynamic scenes: MLP, tensor factorization, and voxel representations. D-NeRF [36] leverages a large deformation MLP to model the transition and the trajectories of the objects and extends NeRF to simulated dynamic scenes. DyNeRF [28] uses long latent vectors to encode temporal information along with large MLPs to achieve high-fidelity novel view synthesis in real dynamic scenes. Park et al. [34] propose an interpolation-based MLP representation for dynamic scenes and achieve significant acceleration over DyNeRF. HyperReel [1] proposes a network to predict the sample locations and represent dynamic scenes by key-frame-based TensoRF [9] representation to ensure high fidelity and real-time rendering without any custom CUDA kernels. K-Planes [17] and HexPlane [7] factorize the whole 4D field to the product of 2D planes. Though differently implemented, they both achieve high-fidelity results with intermediate model size (200MB) and training time (100 minutes) on DyNeRF dataset. TineuVox [16] leverages the deformation MLP in D-NeRF and implements a 3D voxel grid with multi-scale interpolation. It achieves high fidelity and fast convergence on D-NeRF dataset. MixVoxels [45] leverages the standard deviation across the time of all pixels to estimate a variation field to decompose the dynamic scene into static and dynamic voxels. Equipped with different representations for static and dynamic objects, MixVoxels achieves high-fidelity results and super-fast convergence (15 minutes) on DyNeRF dataset. Song et al. [43] train Instant-NGP [31] and TensoRF [9] offline and stream the models along the feature channels for dynamic scenes. Concurrent works are leveraging 4D Gaussian point clouds [54, 49] or 3D Gaussians with deformation networks [55] achieving real-time rendering and high fidelity in dynamic scenes.
2.3 Continual learning of NeRF
Continual learning (online learning) aims to adapt to new tasks while defying catastrophic forgetting (resisting performance downgrade on the previous tasks) [12]. There are mainly 3 approaches to achieving continual learning: replay, regularization, and parameter isolation. The replay-based methods either store the previous data in a buffer and replay them [39], or use the model to generate the pseudo-ground truth of previous tasks while training on new tasks [41]. The regularization-based methods hold the parameters close to the solution of previous tasks [14]. The parameter-isolation methods use different parameters to cope with different tasks. Recent efforts based on generative replay [6, 35, 11] enable continual learning of NeRF in static scenes and achieve high fidelity. StreamRF [27] applies continual NeRF to dynamic scenes by parameter isolation. Their explicit voxel representation enables fast reconstruction (75 minutes for 300 frames) and could be streamed frame by frame. INV [48] uses a shared color MLP and trains different structure MLPs frame by frame to form a continual NeRF representation and achieves high-fidelity results in dynamic scenes.
3 Methodlogy
3.1 Preliminaries
Inspired by the previous NeRF variants [36, 28, 17, 7, 45, 31, 38] and continual learning methods [6, 35, 11, 27, 43, 48], we focus on the dynamic novel view synthesis problem in the multi-view setting and solve the problem in a continual learning manner to alleviate the huge memory consumption and bandwidth costs.
Following previous methods [36, 28, 17, 7], we consider the dynamic scene as a set of 3D points $\mathbf{p}=(x,y,z)$ with optical attributes $(\sigma,\mathbf{c})$, where $\sigma$ is the volume density of the point and $\mathbf{c}$ is the color of the point. We train our model along camera rays to reconstruct the scene. A camera ray is defined as:
$\mathbf{r}(u)=\mathbf{o}+u\mathbf{d},$ (1)
where $\mathbf{o}$ is the optical center of the camera, $\mathbf{d}$ is the direction off the camera. For each point along the camera ray with coordinates $(x,y,z)$ observed at direction $\mathbf{d}=(\phi,\psi)$, a field function $F$ is queried to obtain the optical attributes $(\sigma,\mathbf{c})$. Conditioned on time $t$, the entire model is defined as:
$(\sigma,\mathbf{c})=F_{\Theta}(x,y,z,t,\phi,\psi),$ (2)
where the function $F$ denotes the dynamic NeRF model and $\Theta$ denotes trainable model parameters.
Let $I_{i}$ denote the $i^{th}$ input image corresponding to observation direction $d_{i}$, $\mathbf{V}=\{\nu_{i}~{}|~{}\nu_{i}=(I_{i},d_{i}),1\leq i\leq N_{view}\}$ is the set of the training views, where $i,N_{view}\in\mathbb{Z^{+}}$ and $\mathbf{D_{a}}$ is the readily-accessible data at one step. Previous methods [6, 35] for static scenes define the continual learning problem as learning a unified representation with a limited subset of views $V_{a}$ available at every optimization step, i.e. $\mathbf{D_{a}}=V_{a}\subsetneq V$. However, a more natural definition of continual NeRF is in the dynamic scenes, where the observations at different time stamps arrive continually. Denoting all the frame times as $\mathbf{T}=\{\tau_{i}~{}|~{}1\leq i\leq N_{time}\},i,N_{time}\in\mathbb{Z^{+}}$, we define the continual learning problem as synthesizing a video sequence when the model could only have access to a limited number of frames $\mathbf{T_{a}}\subsetneq\mathbf{T}$ from all views $\mathbf{V}$ at each step, i.e. $\mathbf{D_{a}}=\{\mathbf{V}\times\mathbf{T_{a}}\}$ where $\times$ is the Cartesian product of sets.
To simplify the notations, let the whole sequence of frames be evenly divided into $k$ chunks where a chunk $\mathbf{T_{k}}=\{\tau_{i}~{}|~{}\lfloor\frac{N_{time}}{T_{chunk}}\rfloor=k\}$ is a sequential subset of length $T_{chunk}$. We define the continual learning problem as learning a representation where there are only the frames in the current chunk available at each optimization step, i.e. $\mathbf{T_{a}}=\mathbf{T_{k}}$. Accordingly, $T_{chunk}=1$ denotes training the model frame by frame, $T_{chunk}=10$ denotes training the model with a buffer of 10 frames, $T_{chunk}=300$ denotes training the entire 300-frame sequence offline. The definition is close to real-world applications since caching all training frames in memory or storing them on the disk is limited in scalability for long videos.
3.2 Representation of Short Dynamic Intervals
As pointed out by [46], naive 4D hash grids are not suitable for high-fidelity dynamic view synthesis. To achieve fast reconstruction and slim model size simultaneously, we represent time and spatial features separately and concatenate them to alleviate feature confusion. The spatial hash features may be encoded along 3D voxels [31], 2D planes [17, 21], or both [38]. The time feature is encoded per frame following [28]. Following previous state-of-the-art methods [30, 31, 16], we apply a two-fold field function for geometry and color individually. The architecture of the field function is thus defined as:
$(\sigma,\mathbf{H})=f_{\theta_{1}}(\Psi(\mathbf{x})\oplus\Gamma(t)),$ (3)
$\mathbf{c}=f_{\theta_{2}}(\mathbf{H},\phi,\psi).$ (4)
where $\Psi(\mathbf{x})$ denotes configurable spatial hash encodings, $f_{{\theta_{1}}}$ is a tiny MLP with trainable parameter ${\theta_{1}}$, $\Gamma(t)$ is the hash encoding for time axis alone, $\oplus$ is the concatenation of vectors. The MLP $f_{{\theta_{1}}}$ in Eq. 3 fuses features across the spatial domain and time domain and obtains the volume density $\sigma$. View-dependent colors $\mathbf{c}$ are regressed by another tiny MLP with parameter $\theta_{2}$ in Eq. 4. The spatial encoders $\Psi(x)$ follow a multi-resolution concatenated format, defined as:
$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}~{}P_{m}(x,y,z)$ (5)
where $P_{m}(x,y,z)\in\mathbf{R^{F}}$ denotes the tri-linearly interpolated value from the nearby 8 grid vertices w.r.t. 3D coordinates $(x,y,z)$ at the $l^{th}$ resolution in Instant-NGP [31], $L$ is the number of the resolutions. $\Psi(x)$ may also be configured as the fusion of 2D and 3D features in MERF [38], defined as:
$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}~{}P_{m}(x,y,z)+Q_{mxy}(x,y)+Q_{myz}(y,z)+Q_{mzx}(z,x)$ (6)
where $Q_{mij}(i,j)$ denotes the bi-linearly interpolated value from the 4 nearby vertices in the orthogonally projected planes w.r.t. axes $(i,j)$ at the $m^{th}$ resolution. Removing the $P_{m}(x,y,z)$ term leads to a pure plane representation like [17, 21].
3.3 CD-NGP
Unlike the existing replay-based methods for static scenarios [6, 35], we solve the continual learning problem of dynamic novel view synthesis via parameter isolation. The key insight is that the current state-of-the-art methods by tensor factorization [7, 17] or explicit voxels [45] are not accurate estimations of the object trajectory and inherent dynamics, but only simple regressions of the optical attributes in 3D space within a limited time. Therefore, the replay method is of limited use because the new transient processes are constantly arriving and the model of a limited size is not capable of reconstructing unlimited transient effects and thus suffers from catastrophic forgetting.
To address this issue, we propose our Continual Dynamic neural graphics primitives (CD-NGP) based on parameter isolation. We split the model into $N_{chunk}$ branches to represent different chunks of the dynamic scene accordingly. The $k^{th}$ branch contains spatial hash encoder $\Psi_{k}(\mathbf{x})$, temporal hash encoder $\Gamma_{k}(t)$, and MLPs $f_{\theta_{1k}},f_{\theta_{2k}}$. We train the first branch (base branch) with a large spatial hash table to better reconstruct the scenes with initial movements and static objects. Then we reuse the spatial features from the base branch and train the subsequent branches (auxiliary branches) to compensate for residuals. We thus substitute the $\Psi(\mathbf{x})$ in Eq. 3 with the fused feature to obtain the volume density $\sigma$ at the $k^{th}$ model branch under the continual learning setting as:
$(\sigma,\mathbf{H})=f_{\theta_{1k}}((\Psi_{0}(\mathbf{x})+\Psi_{k}(\mathbf{x}))\oplus\Gamma_{k}(t)),$ (7)
where $\Psi_{0}(\mathbf{x})$ denotes the spatial encoder of the base branch, $\Psi_{k}(\mathbf{x})$ denotes the spatial encoder of the current branch. The view-dependent colors are regressed as Eq. 4 with the color MLP $f_{\theta_{2k}}$.
The scalability of our method derives from the redundancy of dynamic scenes and the flexibility of hash table sizes. As pointed out by [45], most objects ($>90\%$) in the dynamic scenes are static. Therefore, the first chunk could be represented by a large hash table to reconstruct the scenes with initial movements and static objects, the subsequent chunks could be represented by small hash tables to compensate for errors, and the whole representation could be designed to leverage the fusion of features from the corresponding two branches to ensure scalability. To be specific, if we follow the default setting of Instant-NGP [31] and set the hash table in the first branch as $2^{P_{1}}=2^{19}$, the sizes of hash tables in subsequent branches could be configured as $2^{P_{2}}=2^{14}$, which consumes only $1/32$ of the parameters in the first branch. Since the spatial hash tables consume the most parameters in the model, and most chunks are encoded by the subsequent branches, the advantage of our asymmetric hash table structure is obvious: compared with the previous linear-streaming method [43], the total bandwidth consumption is reduced from $\mathcal{O}(\mathbf{K}\cdot 2^{19}LF)$ to $\mathcal{O}(N_{chunk}\cdot 2^{14}LF+2^{19}LF)$, where $\mathbf{K}$ is their bandwidth budget controlling the bitrates in streaming scenarios. If the bandwidth is extremely limited, the hash tables in our model could be arranged as Eq. 6 to reduce the bandwidth cost additionally.
3.4 Rendering and Optimization
Given the sampled optical attributes on the rays in Eq. 1, we adopt volume rendering to get the 2D pixels, defined as:
$\hat{C}(r)$ $=\sum_{i=1}^{N}T_{i}\alpha_{i}\mathbf{c_{i}}=\sum_{i=1}^{N}T_{i}(1-e^{\sigma_{i}\delta_{i}})\mathbf{c_{i}},$ (8)
where $T_{i}=\exp(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j})$ denotes the transmittance, $\delta_{i}$ is the interval between neighboring samples, $N$ is the number of samples along the ray. The model is optimized by L2 photometric loss with three regularization terms, defined as:
$L=\frac{1}{|\mathbf{R}|}\sum_{r\in\mathbf{R}}\|C(r)-\hat{C}(r)\|_{2}^{2}+\lambda_{d}L_{d}+\lambda_{o}L_{o}+\lambda_{r}L_{r}$ (9)
where $\mathbf{R}$ denotes the set of the camera rays. $L_{d}$ is the distortion loss in Mip-NeRF 360 [3], defined as:
$L_{d}$ $=\sum_{i,j}w_{i}w_{j}\left\lvert\frac{s_{i}+s_{i+1}}{2}-\frac{s_{j}+s_{j+1}}{2}\right\rvert$
$+\frac{1}{3}\sum_{i}w_{i}^{2}(s_{i+1}-s_{i})$ (10)
where $s_{i}$ denotes the normalized ray distance along the ray, $w_{i}=T_{i}\alpha_{i}$ is the weight of the $i^{th}$ sample. $L_{o}$ is the entropy on the opacity by [25], defined as:
$L_{o}=-o\log(o),\quad o=\sum_{j=1}^{N}T_{j}\alpha_{j}.$ (11)
$L_{r}$ is L1 regularization on the spatial features in the current branch as:
$L_{r}=||\Psi_{k}(\mathbf{x})||_{1}.$ (12)
We set $\lambda_{r}$ as 0.001 to encourage leverage of the base hash encoder. Based on [25], we set a larger $\lambda_{d},\lambda_{o}$ as 0.005 for alleviating foggy effects.
4 Experiments
4.1 Datasets
DyNeRF dataset [28] is the most prevailing dataset for dynamic view synthesis that provides high-quality multi-camera videos, on which most of the current state-of-the-art methods are evaluated. DyNeRF dataset contains 6 individual cooking scenes with varied illumination, food topological changes, and transient effects like flames. It contains five 300-frame videos and a 1200-frame video, with up to 21 synchronized camera views at the resolution of $2704\times 2028$ at 30 frames per second (FPS). To further prove the viability of our method on long video sequences, we provide a new dataset containing 5 dynamic scenes at $3840\times 2160$ resolution, the lengths of which are at least 200 seconds. Similar to the DyNeRF dataset [28], scenes in our dataset are recorded by 20 GoPro cameras at 30 FPS. Compared with those in the DyNeRF dataset, the dynamics are much harder, including large motions of human bodies and transient effects in the background. Please refer to the supplementary material for more details.
Figure 2: Example frames sampled from the proposed long multi-view video dataset.
4.2 Implementation details
Our implementation is based on HexPlane [7], Instant-NGP [31], and NGP-pl [25]. All the metrics are measured on a single RTX 3090 GPU if not otherwise specified. Similar to the previous work [7, 28], we select the view in the center with index 00 for evaluation on both datasets. Following [7], we exclude the unsynchronized coffee martini scene and conduct experiments on the others. Following most of the state-of-the-art methods [28, 27, 17, 45, 43] we train and evaluate the model at resolution $1352\times 1014$ on the DyNeRF dataset. We use $1280\times 720$ resolution for our dataset to fit into memory. We use the importance sampling scheme in [28] at a batch size of 1024 rays. The parameters are optimized by Adam [23], with $\epsilon=10^{-15},\beta=(0.9,0.96)$. The learning rate is set as 0.001 and scheduled by cosine function following [25]. We set the size of the hash table in the first branch as $2^{19}$ and the size of the hash table in the subsequent branches as $2^{14}$, and we set the structural parameters $(L,F)$ in the spatial hash encoders as $(12,2)$ if not otherwise specified. At the base branch, we initialize the models and train for 18k iterations, and for each of the subsequent branches we train for 3k iterations using the previous branch for initialization. We use LeakyReLU activation in the MLP regressors. For long video sequences, we assign 30 chunks as an episiode. The MLP regressors are re-trained from initialization in the first branch of each episode to adapt to recent changes. Strategies for updating the occupancy grids are identical with [31]. To save storage, we use only one shared occupancy grid for all the branches. We report the model sizes with the occupancy grid under bfloat16 precision and parameters in float32 precision. Please see the supplementary material for more details.
4.3 Results and Comparisons
Table 1: Comparisons between our CD-NGP and SOTA methods on the DyNeRF dataset. Memory denotes the maximum host memory required for training. - denotes the metrics unreported in the original publications. denotes the methods measured only on the most challenging flame salmon scene. Best, Second, and Third metrics are in highlight.
[…]
offline DyNeRF [28] 300 29.58 0.020 0.083 7 days 28MB >100GB
NeRF-T [28] 300 28.45 0.023 0.10 - - >100GB
Interp-NN [34] 300 29.88 - 0.096 2 days 20MB >100GB
HexPlane [7] 300 31.57 0.016 0.089 96 min 200MB 62GB
K-Planes [17] 300 31.63 - - 110 min 200MB >100GB
MixVoxels-T [45] 300 30.71 0.024 0.162 15 min 500MB >100GB
4DGS [54] 300 32.01 - 0.055 6 hours 4000MB <10GB
HyperReel [1] 50 31.1 - 0.096 9 hours 360MB 18GB
NeRFplayer [43] 300 30.29 - 0.152 5.5 hours - >100GB
online StreamRF [27] 1 28.26 0.029 0.226 75 min 5000MB <10GB
HexPlane-PR 10 24.03 0.081 0.244 100 min 318MB 14GB
HexPlane-GR 10 25.17 0.050 0.272 125 min 318MB 14GB
INV [48] 1 28.72 - 0.121 40 hours 336MB <10GB
CD-NGP (Ours) 10 30.23 0.029 0.198 75 min 113MB 14GB
4.3.1 Standard-length sequences.
Table 1 shows the quantitative comparisons between CD-NGP and other state-of-the-art methods on DyNeRF dataset. We use the officially released metrics. [28, 34] are trained on 8 V100 GPUs or equally 4 A100 GPUs. [28, 34, 43] have no official implementations available. We are unable to obtain the missing quality metrics in [17] because of their huge memory occupation in the default resolution. [7] uses resolution $(1024,768)$ in their implementation, so the memory occupation is much smaller than other offline methods. We use the released output of [27] for DSSIM and LPIPS metrics. [27, 48] are methods trained per-frame. [43] is trained on the whole video sequence but streamable upon inference. 4DGS [54] caches the frames onto disk for training to alleviate memory occupation with a trade-off on training speed. The reconstruction quality is measured by PSNR, DSSIM, and LPIPS [57] (by AlexNet [24]) following common practices [28, 7, 45]. To prove the viability of our parameter-isolation-based method, we apply the previous state-of-the-art method HexPlane [7] to continual learning setting with parameter-regularization (HexPlane-PR) and generative replay (HexPlane-GR) as additional baselines.
(a) HexPlane-offline
(b) MixVoxels-T
(c) StreamRF
(d) HexPlane-PR
(e) HexPlane-GR
(f) CD-NGP (ours)
Figure 3: Comparison of reconstruction quality on the most challenging flame salmon scene in DyNeRF dataset. PSNRs are computed under the resolution $1352\times 1014$.
We compare our method with both online and offline baselines. As demonstrated in Tab. 1, our CD-NGP is 32×32$\times$32 × faster than the MLP-based continual representation [48], $\times4$ smaller than the voxel-based continual representation [27] and significantly outperforms the two additional baselines in the continual learning setting. The undesirable results from HexPlane-PR and HexPlane-GR show the limitations of the regularization-based and replay-based methods in dynamic scenes: unlike the static scenes, the objects are not stationary across time and the models with limited complexity are not capable of reconstructing unlimited transient effects. Benefiting from the parameter isolation like previous methods [27, 48], our method achieves fast convergence and compact model size simultaneously: it leverages the hash tables to ensure fast convergence and reuses the features from differently-configured spatial hash tables to ensure high scalability. Compared with prevailing offline methods [28, 34, 7, 17, 45, 54], our method only requires caching $T_{chunk}=10$ frames into buffer by continual learning and thus reduces memory occupation by $85\%$. Our method also achieves $40\%$ less model size ($0.4$MB/frame) by leveraging the feature fusion of the base and the auxiliary branches with comparable quality.
Figure 3 provides qualitative comparisons. Compared with offline methods in Fig. 3(a), Fig. 3(b) and the online method in Fig. 3(c), our model is much smaller without large quality loss. Compared with regularization-based and replay-based continual learning methods in Fig. 3(d) and Fig. 3(e) respectively, our method achieves significantly higher quality. Results of other scenes including depth maps are shown in Fig. 4. As illustrated in Fig. 6, our method reaches a new balance between rendering quality, model size, and training time while remaining memory-friendly, in which we use metrics on the flame salmon scene for fair comparison. The perceptual consistency is measured by $1-$ LPIPS, and the speed is the reciprocal of the training time. The compactness and RAM-Efficiency are mapped by $-log$ function from the model size and memory occupation respectively.
Figure 4: The results of the proposed CD-NGP on DyNeRF dataset.
4.3.2 Long sequences.
We also provide quantitative comparisons on our proposed long multi-view video dataset in Tab. 3 and Fig. 6 to further show the advantage in scalability of our method. We split the sequence into 300-frame chunks for the compared methods since the offline baseline [7] is designed for 300-frame sequences in the DyNeRF dataset. We set the hash table size as $2^{12}$ in the auxiliary branches and use $T_{chunk}=5$ in our model. The offline baseline [7] produces unstable results on one of the 5 scenes in our dataset, so we only use the other 4 scenes to obtain the metrics of [7]. As shown in Tab. 3, our method reduces memory and storage consumption by at least 4×4$\times$4 × with comparable quality to the offline baseline. Compared with the results in Tab. 1, the performance gap in storage consumption is even larger in long video sequences, because the auxiliary branches consume the majority of the parameters and the complexity in storage is reduced to $\mathcal{O}(N_{chunk}\cdot 2^{12}LF)$. The PSNR curves in Fig. 6 further demonstrate the stability of our method. The offline baseline produces the best quality. The compared online methods suffer from catastrophic forgetting: the PSNRs at the beginning of the chunks are much lower than at the end. Our method resists the problem by parameter isolation and ensures the quality of the whole sequence. We also provide comparisons between rendered results in Fig. 8. The regularization method in Fig. 7(a) reconstructs the TV screen in the background erroneously. The replay method in Fig. 7(b) fails to recover the details. Ours in Fig. 7(d) produces comparable quality to the offline baseline in Fig. 7(c).
Figure 5: Comparison between different performance indexes of our method and recent representative baselines.
Figure 6: PSNR curves of the seminar scene. The solid line and dotted line denotes offline and online methods respectively.
4.4 Ablation studies
(a) HexPlane-PR
(b) HexPlane-GR
(c) HexPlane
(d) CD-NGP (ours)
Figure 7: Comparison of reconstruction quality on the 19th frame of the seminar scene in our long video dataset.
Figure 8: PSNR curves of asymmetric (solid line, ours) and symmetric (dotted line) hash table designs on the sear steak scene.
4.4.1 Lengths of hash tables.
Benefiting from the multi-branch hash table design, our model is highly efficient and configurable. We change the parameter numbers of the initial branch and the auxiliary branches and obtain the results on the DyNeRF dataset in Tab. 3. $(P_{1},P_{2})$ are the log2 values of the hash table lengths of the base branch and the other branches. Please note the whole model sizes include the sizes of the occupancy grid and are thus larger than those deduced from parameter numbers. Larger hash tables alleviate hash collisions and lead to better results. However, the default configuration $(19,14)$ in CD-NGP produces better results than the naive $(16,16)$ configuration even if the storage consumption is $40\%$ lower. Moreover, the $(19,12)$ configuration of CD-NGP significantly outperforms the naive $(14,14)$ configuration at a similar storage cost. To further demonstrate the scalability of our method, we provide the PSNR curves of different hash table designs on the sear steak scene in Fig. 8. Equipped with a larger hash table in the base branch, the asymmetric setting $(19,14)$ and $(19,12)$ outperforms others in the beginning and produces comparable metrics to the naive $(16,16)$, $(14,14)$ configuration respectively in the subsequent chunks. The advantage comes from our asymmetric hash table designs: the base branch with a large hash table recovers more environment patterns, and the auxiliary branches could thus use fewer parameters to cope with recent changes only.
Table 2: Comparisons between our CD-NGP and baselines on the long video sequences in our dataset. denotes the method trained offline. Metrics are measured on the first 1200 frames. Compared baselines are trained with a chunk size of 300 frames.
[…]
HexPlane-PR 20.90 0.286 1272MB 12GB
HexPlane-GR 24.82 0.282 1272MB 12GB
CD-NGP (ours) 26.50 0.201 201MB 12GB
HexPlane [7] 28.32 0.117 800MB 74GB
Table 3: Ablation study on hash table sizes. denotes the default setting.
[…]
$(18,18)$ 151M 30.55 638MB
$(16,16)$ 41M 30.08 191MB
$(14,14)$ 12M 29.32 75MB
$(19,14)^{{\dagger}}$ 21M 30.23 113MB
$(16,14)$ 13M 29.78 79MB
$(19,12)$ 14M 30.09 82MB
4.4.2 Spatial representations.
Table 4: Ablation study on spatial representations. Different representations lead to subtle changes in quality metrics.
[…]
CD-NGP 21M 30.23 0.198 75min 113MB
Voxel-C 21M 30.52 0.174 80min 114MB
Plane 14M 29.16 0.215 77min 83MB
MERF 7.5M 29.93 0.201 109min 57MB
MERF-C 7.5M 30.22 0.174 115min 57MB
Table 5: Ablation study on regularization weight $\lambda_{r}$. $10^{-3}$ is the default value.
[…]
$10^{-5}$ 30.20 0.0269 0.198
$10^{-4}$ 30.18 0.0268 0.196
$10^{-3}$ 30.23 0.0269 0.198
$10^{-2}$ 30.09 0.0279 0.197
1 29.66 0.0313 0.205
The proposed hash-based pipeline is not only practical with voxel hash grids but also compatible with other hash-based spatial backbones. We provide quantitative results on plane representation like [17, 21] and hybrid representation (MERF) [38] on the DyNeRF dataset in Tab. 4. Since the element-wise product for plane representation in [17] produces unsatisfactory results due to the hash collision, we remove the $P_{m}(x,y,z)$ in Eq. 6 and fuse the features $Q_{mij}(i,j)$ by concatenation in the implementation.
Table 4 also shows metrics of different feature fusion methods. We replace the element-wise sum in Eqs. 7 and 6 with vector concatenation and denote them as Voxel-C and MERF-C in Tab. 4. The hybrid representations (MERF and MERF-C) achieve comparable model sizes ($\sim$60MB) to methods for static scenarios [31, 21] without significant quality loss, reducing the average bandwidth cost to $0.2$MB/frame. Benefiting from larger feature dimensions, the vector concatenation leads to slightly better quality (especially when measured in LPIPS) by more training time, and the reasonable trade-off also shows the robustness of our method. Compared with the hash table designs in Tab. 3, the Voxel-C representation in Tab. 4 almost performs as well as the bulky $(18,18)$ configuration in Tab. 3 and reduces the model size by $80\%$.
We also provide ablation studies on the regularization weights of auxiliary spatial encodings. We apply regularizations on $\Psi_{k}(x)$ with different $\lambda_{r}$ and show the results in Tab. 5. The regularization is of slight improvement, and our method is robust within a wide range of $\lambda_{r}$. Please refer to the supplementary material for more ablation studies.
5 Conclusion and Limitation
We propose CD-NGP, a continual learning framework for multi-view dynamic scenes based on parameter isolation. Equipped with the multi-branch hash encoding structure, it shows high efficiency and scalability and is capable of reconstructing long video sequences with low host memory occupation. Compared with offline methods, it consumes $40\%$ less bandwidth and $85\%$ less memory than recent explicit baselines and achieves comparable quality while maintaining the speed advantage over the pure-MLP-based methods. Compared with online baselines, it significantly outperforms replay and regularization-based continual learning methods and reaches a new balance among reconstruction quality, training speed, bandwidth occupation, and memory cost. However, like most of the NeRF-based methods, it still requires querying the field function and is not appropriate for achieving real-time rendering. Further efforts could be made to apply baking methods or Gaussian point clouds to continual learning to achieve real-time rendering in long video sequences for wide applications.
Appendix 0.A Detailed Results on DyNeRF Dataset
We list the per-scene results of the proposed continual learning method with different spatial representations on the DyNeRF dataset in Tab. 6.
Table 6: Detailed results of the proposed CD-NGP and other hash-based methods on the DyNeRF dataset.
Model Flame Salmon Cook Spinach
[…]
CD-NGP 26.37 0.04586 0.2886 30.76 0.02440 0.1793
Voxel-C 26.65 0.04243 0.2680 31.03 0.02288 0.1593
Plane 25.48 0.06058 0.3173 29.41 0.03487 0.2010
MERF 26.15 0.04952 0.2921 30.35 0.02931 0.1858
MERF-C 26.56 0.04435 0.2546 30.69 0.02609 0.1583
Model Flame Steak Sear Steak
[…]
CD-NGP 31.27 0.02132 0.1736 31.71 0.01917 0.1701
Voxel-C 31.64 0.01823 0.1403 32.02 0.01698 0.1424
Plane 30.26 0.02949 0.1804 30.56 0.02723 0.1811
MERF 31.01 0.02508 0.1719 31.44 0.02181 0.1698
MERF-C 31.13 0.02424 0.1506 31.66 0.01989 0.1413
Model Cut Roasted Beef Average
[…]
CD-NGP 31.03 0.02366 0.1801 30.228 0.026882 0.19834
Voxel-C 31.26 0.02205 0.1578 30.520 0.024514 0.17356
Plane 30.11 0.03302 0.1951 29.164 0.037038 0.21498
MERF 30.71 0.02765 0.1856 29.932 0.030674 0.20104
MERF-C 31.04 0.02450 0.1632 30.216 0.027814 0.17360
Appendix 0.B More Implementation Details
0.B.1 Parameter numbers and streaming efficiency.
We set the dimensions of the latent feature $\mathbf{H}$ in Eq. 3 as 48. We use MLPs with two hidden layers and the dimensions of the hidden layers in Eqs. 3 and 4 are set as 128, and 64 respectively. We use MLPs with identical structures for different branches. We set the resolutions of the hash encoders of the proposed CD-NGP and MERF representations within $(16,2048)$ times the scale of the scene, and we use $(64,2048)$ for the plane representation.
Following MERF [38] for static scenes, we use a smaller hash table for the MERF and the MERF-C than pure voxel hash grid [31] representation in Tab. 4 to ensure the compactness and scalability of our method. Denoting the lengths of hash tables along pure voxels in a branch of CD-NGP as $2^{P}$, the lengths of hash tables in the MERF and the MERF-C are set as $2^{P-3}$ and $2^{P-4}$ for voxel and plane encoders respectively. The number of the parameters of the MERF representation is thus reduced to $1-2^{-3}-3\times 2^{-4}=31.25\%$ of the voxel hash grids. Since the hash tables take up the majority of the parameters, the model size is reduced from $21M$ to $7.5M$ as in Tab. 4. We use $2^{P-2}$-length hash tables for the Plane representation in Tab. 4 to balance the quality and the model size. The structural configurations $(L,F)$ of the plane representation in Tab. 4 are set as $(6,4)$ for better results.
The detailed parameter numbers and bandwidth consumption are thus listed in Tab. 7. 3D Enc. and 2D Enc. denote the spatial encodings along voxels and planes respectively. The tuples $(M_{1},M_{2})$ in the cells denote the parameter numbers of the base encoder and each of the auxiliary encoders. Following [27, 43, 1, 48] we also report the minimum per-frame bandwidth consumption $\mathbf{B_{min}}$, which only depends on the parameter numbers of the auxiliary branches and is thus significantly lower than the average bandwidth consumption $\mathbf{B_{avg}}$. Since the base encoder and the density grid could be shared among different branches and transmitted only once in the beginning, the minimum bandwidth consumption is reduced to $0.065\sim 0.16$ MB per frame, which is 100×100$\times$100 × less than the voxel-based StreamRF [27] and 7×7$\times$7 × less than the pure MLP representation INV [48].
Table 7: The parameter numbers of each module of the 30-branch representations in Tab. 4. $\mathbf{B_{min}}$ and $\mathbf{B_{avg}}$ are minimum and average bandwidth costs in MB/frame. ##\## denotes the number of parameters per branch (10 frames).
Model ##\##3D Enc. ##\##2D Enc. ##\##MLP ##\##Online ##\##Total $\mathbf{B_{min}}$ $\mathbf{B_{avg}}$
CD-NGP $(9.9M,0.37M)$ $0.042M$ $0.39M$ $21M$ 0.16 0.38
Voxel-C $(9.9M,0.37M)$ $0.046M$ $0.40M$ $21M$ 0.16 0.38
Plane - $(1.9M,0.082M)$ $0.048M$ $0.28M$ $14M$ 0.11 0.28
MERF $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.042M$ $0.16M$ $7.5M$ 0.061 0.19
MERF-C $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.052M$ $0.17M$ $7.5M$ 0.065 0.19
0.B.2 Continual learning pipeline.
We further illustrate the whole process of our method in Algorithm 1, where $\eta_{init}$ and $\eta_{aux}$ denote the numbers of iterations of the initial branch and the auxiliary branches, respectively. $\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)}$ denote the model parameters in the branches. The whole model is trained chunk by chunk. The implementation will be available upon publication.
Algorithm 1 Continual learning pipeline
0:  $N_{chunk},T_{chunk},\eta_{init},\eta_{aux},\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)},i\in\mathbb{N}$
0:  The whole representation for dynamic scenes.
[…]
2:  while $N_{chunk}$ do
[…]
5:     if $k~{}\mathbf{mod}~{}T_{episode}~{}=0$  then
[…]
7:     else
[…]
10:     end if
11:     Train the $k^{th}$ model branch $(\Psi_{k},\Gamma_{k},\theta_{1}^{(k)},\theta_{2}^{(k)})$ for $\eta$ steps.
[…]
13:  end while
Appendix 0.C More Details about the Long Video Dataset
We collect the dataset for evaluating the scalability of our method and further research. We have checked the videos to make sure they do not contain controversial or offensive content before using them for research purposes. The faces in the videos may contain identifiable information, but the information will neither be used for any other purposes nor affect the anonymity of the submission. All videos are collected with the explicit consent of all the participants. The view in the upper center is selected for validation. We synchronize the videos by environmental sound, and we put chessboard images in the scene to help calibrate the cameras and estimate the camera poses.
Appendix 0.D More Ablation Studies
0.D.1 Temporal representations.
Our continual learning pipeline is also compatible with a variety of representations for the time axis. Currently, there are 2 other prevailing representations for the time axis alone: pure frequency encoding [36] and MLP with frequency-encoding [16]. We replace the temporal hash encoding in our method with them on DyNeRF dataset and show the results in Tab. 9. Our time latent code slightly outperforms others on DyNeRF dataset. The robustness against changes in the temporal representations also shows the viability of the continual learning method.
Table 8: Ablation study on temporal representations.
[…]
CD-NGP 30.23 0.0269 0.198
Time-MLP 30.20 0.0273 0.202
Time-Freq 30.19 0.0272 0.203
Table 9: Ablation study on field composition methods.
[…]
CD-NGP 30.23 0.0269 0.198
Full composition 29.24 0.0367 0.228
Static-dynamic 28.27 0.0568 0.291
0.D.2 Field composition methods.
We also conduct ablation studies on the field composition methods and show the results in Tabs. 9 and 9. The field composition process is defined as:
$(\sigma,\mathbf{c})=(\sigma_{i}+\sigma_{a},\mathbf{c}_{i}+\mathbf{c}_{a}).$ (13)
Full composition in Tab. 9 denotes the naive composition of 2 fields from the base branch and the current branch respectively, i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ are the field from the base branch and $\sigma_{a}$ and $\mathbf{c}_{a}$ in Eq. 13 are the field from the current branch. Both branches are encoded with the same architecture as Eq. 3. Static-dynamic in Tab. 9 denotes a naive composition of the field from a static branch by the base encoder and a dynamic branch by the auxiliary encoders similar to [45], i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ in Eq. 13 are encoded without the $\Gamma(t)$ in Eq. 3. The online training strategies remain the same. Our CD-NGP outperforms the other methods in all metrics. Furthermore, as compared in Fig. 9, the compared composition methods in Figs. 9(c) and 9(d) suffer from salt and pepper noise because of the explicit composition, while our CD-NGP circumvents the problem by fusing features in the latent space and produces better results.
(a) Ground truth
(b) CD-NGP (Ours)
(c) Full composition
(d) Static-dynamic
Figure 9: Comparison of different field composition methods.
0.D.3 Chunk size.
We also conduct ablation studies on the chunk sizes and show the results in Tabs. 11 and 11. We only report the results on the cook spinach scene in the DyNeRF dataset for brevity. Our method is robust within a wide range of chunk sizes. As in Tab. 11, the smaller chunk size leads to better results when controlling the iteration numbers of each chunk. When the total iterations of the whole 300-frame sequence are held constant, the larger chunk size leads to better results as in Tab. 11. We do not cherry-pick the best chunk size and select $T_{chunk}=10$ as the default configuration in our method. The selection of the chunk size is a trade-off between the quality, robustness, training time, and host memory occupation and should depend on the data.
Table 10: Ablation study on chunk sizes. The iteration numbers of each chunk are held constant under different $T_{chunk}$ settings.
[…]
30 30.31 0.02654 0.1842 80MB
15 30.54 0.02557 0.1816 97MB
10 30.76 0.02440 0.1793 113MB
5 30.99 0.02268 0.1674 162MB
Table 11: Ablation study on chunk sizes. The total iterations of the whole 300-frame sequence are held constant under different $T_{chunk}$ settings.
[…]
30 30.84 0.02392 0.1505 80MB
15 30.76 0.02413 0.1684 97MB
10 30.76 0.02440 0.1793 113MB
5 30.39 0.02602 0.2075 162MB