### v5

# CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes

Abstract
Current methods for novel view synthesis (NVS) in dynamic scenes encounter significant challenges in managing memory consumption, model complexity, training efficiency, and rendering fidelity. Existing offline techniques, while delivering high-quality results, face challenges from substantial memory demands and limited scalability. Conversely, online methods struggle to balance rapid convergence with model compactness. To address these issues, we propose continual dynamic neural graphics primitives (CD-NGP). Our approach leverages a continual learning framework to reduce memory overhead, and it also integrates features from distinct temporal and spatial hash encodings for high rendering quality. Meanwhile, our method employs parameter reuse to achieve high scalability. Additionally, we introduce a novel dataset featuring multi-view, exceptionally long video sequences with substantial rigid and non-rigid motion, which is seldom possessed by existing datasets. We evaluate the reconstruction quality, speed and scalability of our method on both the established public datasets and our exceptionally long video dataset. Notably, our method achieves an $85\%$ reduction in training memory consumption (less than 14GB) compared to offline techniques and significantly lowers streaming bandwidth requirements (less than 0.4MB/frame) relative to other online alternatives. The experimental results on our long video sequences dataset show the superior scalability and reconstruction quality compared to existing state-of-the-art approaches.
keywords:
Continual learning, Dynamic novel view synthesis, Neural radiance field.
[…]
The outline of our continual learning pipeline for dynamic scenes. Our method segments input videos into multiple chunks and encodes the chunks by separate spatial and time hash encodings. Subsequently, we train the auxiliary branches incrementally on each chunk and leverage the re-use of parameters in the base branch to ensure high scalability. The spatial hash features could be encoded according to voxel grids, orthogonally projected plane grids or a combination of both.
Introduction
Neural radiance field (NeRF) introduced in [30] has made significant progress in 3D scene reconstruction and novel view synthesis. It takes multi-view RGB images with their camera poses as input and leverages multi-layer perceptrons (MLPs) to reconstruct both the geometry and color of the scene. Many variants have been proposed to improve NeRF’s efficiency [9, 38, 31, 56, 53], reconstruction quality [2, 3, 4], scalability [47, 50], and pose robustness [5, 26, 13, 52, 32, 51]. Recently, 3D Gaussian Splatting (3DGS) [22] proposes an explicit representation with differentiable rendering, which significantly speeds up the reconstruction and enables real-time novel view synthesis. Despite significant progress having been made in this field, the above-mentioned methods assume that the scene is static and may not work properly in dynamic scenes.
To address the dynamic reconstruction problem, a number of efforts have been made in the offline setting, requiring access to the entire dataset during the training stage. For example, the methods in [36, 28, 34, 33, 15, 29, 44] propose to use MLP-based representation and achieve high-quality results. There are also methods that speed up the optimization by using the tensor factorization ([40, 17, 7]) and voxel representations [16, 45]. The more recent continual-learning-based methods [27, 48] (also mentioned as online methods) can train models frame by frame, and streamable representations [43] enable the model to be loaded on the fly at inference. However, the offline methods require huge host memory e.g., taking more than $100GB$ when reconstructing multi-view dynamic videos of 300 frames, and the online methods are weak in balancing the convergence speed and model compactness. These limitations in scalability makes them impractical for the wide application on long videos.
In this paper, we propose continual dynamic neural graphics primitives (CD-NGP), a fast and scalable representation for reconstructing dynamic scenes, which enables dynamic reconstruction in long videos. As illustrated in CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes, CD-NGP derives from three key ideas: dividing the video into segments, encoding spatial and temporal features separately, and reusing features across different segments to ensure high scalability: Firstly, we partition dynamic videos into multiple segments according to time stamps $t$ and perform reconstruction using distinct model branches to facilitate continual learning. Each branch of our model includes MLP regressors, spatial hash encoding for spatial coordinates $(x,y,z)$, and temporal hash encoding for the timestamp $t$. Secondly, the spatial and temporal features are encoded separately to avoid feature confusion in different axes. For each model branch, we sample 3D points $(x,y,z)$ according to frame pixels and cached occupancy grids like [31], and encode the points into spatial features using spatial hash grids. We also encode temporal features using hash grids for the separated $t$ axis. These spatial and temporal features and the viewing direction $\mathbf{d}$ are concatenated and input into MLP regressors for rendering and optimization. Thirdly, we leverage the redundancy in the dynamic scenes and fuse the spatial features of different branches to yield high scalability. We set a larger hash table size ($2^{19}$) in the initial branch (base branch), and set a smaller size ($2^{14}$) for the other branches (auxiliary branches). This asymetric design facilitates reusing the features trained in the base branch, thereby reducing storage consumption and boosting rendering quality.
Figure 1: Comparison between different performance indexes of our method and recent representative baselines. The perceptual consistency is measured by $1-$ LPIPS by AlexNet [57, 24], and the speed is the reciprocal of the training time. The compactness and RAM-Efficiency are mapped by $-log$ function from the model size and memory occupation respectively.
We validate the effectiveness of our method through comprehensive experimental results on both the widely used public datasets and a self-collected dataset. As illustrated in Fig. 1, our method reaches a new balance between rendering quality, model size, and training time while remaining memory-friendly. Compared with prevailing offline baselines, our method consumes $85\%$ less host memory at the training phase, $40\%$ less model size at the inference phase, and delivers comparable results. Compared with online baselines, our method achieves compact model size, high rendering quality, and fast convergence simultaneously. The contributions of this paper are three-fold as follows:
  • We propose a novel method combining separate spatial and temporal hash encodings, achieving high scalability and rendering quality in dynamic scenes.
  • We propose a fast and highly scalable continual learning framework for dynamic scenes. It enables reconstructing long videos with low memory usage and low storage consumption. Our method also outperforms online alternatives and achieves comparable results to offline methods.
  • We provide a dataset of multi-view exceptionally long video sequences ($>3$ minutes) with large non-rigid motion, which is seldom possessed by existing datasets.
1 Related Work
1.1 Fast novel view synthesis in static scenes
The acceleration of NeRFs falls into two categories: training acceleration and rendering acceleration. Plenoxels [18] uses voxel grids to represent the scene and achieves magnitudes of acceleration over MLP backbones. TensoRF [9] decomposes the 3D scene as the product of tensors and achieves fast novel view synthesis. Plane-based methods [17, 21] project 3D points onto 3 orthogonal 2D planes to ensure sparsity and achieve fast convergence. Instant-NGP [31] leverages an occupancy grid and a multi-resolution hash encoding along with optimized tiny MLPs to enable fast novel view synthesis within a minute. The recent 3D Gaussian method [22] represents the scenes as a union of explicit Gaussian point clouds and achieves high fidelity, fast convergence, and real-time rendering on high-end hardware by eliminating the querying burden of spatial representations. KiloNeRF [37] distills the large MLP in NeRF [30] into tiny MLPs to accelerate rendering by thousands of times. Hedman et al. [20] propose the baking method (converting the MLP representation into explicit meshes) to accelerate the rendering process. They alpha-compose the cached features along the ray and use MLPs only once per ray to overcome the computation load of the per-point MLP querying in NeRF. BakedSDF [56] stores the view-dependent features into Gaussian spheres along the reconstructed surfaces, and it achieves real-time rendering on ordinary laptops. MobileNeRF[10] represents scenes into meshes and enables real-time rendering on mobile devices. Light field methods [42, 8, 19] directly represent the scene as a set of rays to eliminate volume rendering and enable real-time applications on mobile devices.
1.2 Representations of dynamic scenes
There are mainly 3 variants of NeRF for dynamic scenes. The first class is pure MLP-based representation. D-NeRF [36] leverages a large deformation MLP to model the transition and the trajectories of the objects and extends NeRF to simulated dynamic scenes. DyNeRF [28] uses long latent vectors to encode temporal information along with large MLPs to achieve high-fidelity novel view synthesis in real dynamic scenes. Park et al. [34] propose an interpolation-based MLP representation for dynamic scenes and achieve significant acceleration over DyNeRF. The second class is tensor factorization-based representation. HyperReel [1] proposes a network to predict the sample locations and represent dynamic scenes by key-frame-based TensoRF [9] representation to ensure high fidelity and real-time rendering without any custom CUDA kernels. K-Planes [17] and HexPlane [7] factorize the whole 4D field into the product of 2D planes. Though differently implemented, they both achieve high-fidelity results with intermediate model size (200MB) and training time (100 minutes) on the DyNeRF dataset. The third class is the explicit voxel-based representations. TineuVox [16] leverages the deformation MLP in D-NeRF and implements a 3D voxel grid with multi-scale interpolation, and it achieves high fidelity and fast convergence on the D-NeRF dataset. MixVoxels [45] leverages the standard deviation across the time of all pixels to estimate a variation field to decompose the dynamic scene into static and dynamic voxels. Equipped with different representations for static and dynamic objects, MixVoxels achieves high-fidelity results and super-fast convergence (15 minutes) on DyNeRF dataset. Song et al. [43] train Instant-NGP [31] and TensoRF [9] offline and stream the models along the feature channels for dynamic scenes. Concurrent works are leveraging 4D Gaussian point clouds [55, 49] or 3D Gaussians with deformation networks [54] achieving real-time rendering and high fidelity in dynamic scenes.
1.3 Continual learning of NeRF
Continual learning (online learning) aims to adapt to new tasks while defying catastrophic forgetting (resisting performance downgrade on the previous tasks) [12]. There are mainly 3 approaches to achieving continual learning: replay, regularization, and parameter isolation. The replay-based methods either store the previous data in a buffer and replay them [39], or use the model to generate the pseudo-ground truth of previous tasks while training on new tasks [41]. The regularization-based methods hold the parameters close to the solution of previous tasks [14]. The parameter-isolation methods use different parameters to cope with different tasks. Recent efforts based on generative replay [6, 35, 11] enable continual learning of NeRF in static scenes and achieve high fidelity. StreamRF [27] applies continual NeRF to dynamic scenes by parameter isolation. Their explicit voxel representation enables fast reconstruction (75 minutes for 300 frames) and could be streamed frame by frame. INV [48] uses a shared color MLP and trains different structure MLPs frame by frame to form a continual NeRF representation. It achieves high reconstruction quality in dynamic scenes.
2 Method
2.1 Problem definition and preliminaries
[…]
Inspired by the previous NeRF variants [36, 28, 17, 7, 45, 31, 38] for dynamic novel view synthesis and continual learning methods [6, 35, 11, 27, 43, 48] in the static scenes, we focus on the dynamic novel view synthesis problem in a multi-view setting. We solve the problem using continual learning to alleviate the huge memory consumption and bandwidth costs. Unlike the existing replay-based continual learning methods for static scenarios [6, 35], we solve the continual learning problem of dynamic novel view synthesis via parameter isolation. This is because the current offline state-of-the-art methods by tensor factorization [7, 17] or explicit voxels [45] are regressions of the optical attributes in 3D space within a limited time instead of the object trajectory and inherent dynamics. Therefore, the replay method is of limited use in our defined scenario: Since new transient processes are constantly arriving, the model of a limited size is not capable of reconstructing unlimited transient effects. Thus, the replay-based method suffers from catastrophic forgetting. The design of our model follow the parameter isolation approach: Let the time stamps be divided into $N_{chunk}$ chunks, the model consists of $N_{chunk}$ branches to represent different chunks of the dynamic scene accordingly.
Following previous methods [36, 28, 17, 7], we consider the dynamic scene as a set of 3D points $\mathbf{x}=(x,y,z)$ with optical attributes $(\sigma,\mathbf{c})$, where $\sigma$ is the volume density of the point and $\mathbf{c}$ is the color of the point. To reconstruct the scene, we train our model along camera rays which are defined as:
$\mathbf{r}(u)=\mathbf{o}+u\mathbf{d},$ (1)
where $\mathbf{o}$ is the optical center of the camera, $\mathbf{d}$ is the direction from the camera to the scenes. For each point along the camera ray with coordinates $\mathbf{x}=(x,y,z)$ observed at direction $\mathbf{d}=(\phi,\psi)$, a field function $F$ is queried to obtain the optical attributes $(\sigma,\mathbf{c})$. Conditioned on time $t$, the entire model is defined as:
$(\sigma,\mathbf{c})=F_{\Theta}(x,y,z,t,\phi,\psi),$ (2)
where the function $F$ denotes the dynamic NeRF model and $\Theta$ denotes trainable model parameters.
2.2 Representation of short dynamic intervals
To reconstruct the individual chunks of the dynamic scenes, we propose a representation of short dynamic intervals, which serves as the structures of the different branches of our model and handle the dynamic scnes in one chunk efficiently. As shown in [46], naive 4D hash grids for dynamic scenes are short in reconstruction quality. To achieve fast reconstruction and compact model size simultaneously, we represent time and spatial features separately and concatenate them to alleviate feature confusion. The spatial hash features can be encoded along 3D voxels [31], 2D planes [17, 21], or both [38]. The time feature is encoded by short hash tables in a per-frame manner. Following previous state-of-the-art methods [30, 31, 16], we apply a two-fold field function for geometry and color individually. The architecture of the field function is thus defined as:
$(\sigma,\mathbf{H})=f_{\theta_{1}}(\Psi(\mathbf{x})\oplus\Gamma(t)),$ (3)
$\mathbf{c}=f_{\theta_{2}}(\mathbf{H},\phi,\psi),$ (4)
where $\mathbf{x}=(x,y,z)$ is the spatial coordinates, $\Psi(\mathbf{x})$ denotes configurable spatial hash encodings, $f_{{\theta_{1}}}$ is a tiny MLP with trainable parameter ${\theta_{1}}$, $\Gamma(t)$ is the hash encoding for time axis alone, and $\oplus$ is the concatenation of vectors. The MLP $f_{{\theta_{1}}}$ in Eq. 3 fuses features across the spatial domain and time domain and obtains the volume density $\sigma$. View-dependent colors $\mathbf{c}$ are regressed by another tiny MLP with parameter $\theta_{2}$ in Eq. 4. The spatial hash encoders $\Psi(x)$ follow a multi-resolution concatenated format, defined as:
$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}\leavevmode\nobreak\ P_{m}(x,y,z),$ (5)
where the $P_{m}(x,y,z)\in\mathbf{R}^{F}$ denotes the tri-linearly interpolated feature from the nearby 8 grid vertices w.r.t. 3D coordinates $(x,y,z)$ at the $m$-th resolution, $L$ is the number of the resolutions, and $F$ is the dimensions of features $P_{m}(x,y,z)$, which are similar to those in Instant-NGP [31]. The voxel-based hash encoding structure is flexible and could be replaced by other spatial hash encodings, which will be discussed in the experiments.
2.3 Continual dynamic neural graphics primitives (CD-NGP)
As defined in Section 2.1, we split the model into $N_{chunk}$ individual branches to represent the $N_{chunk}$ chunks of the dynamic scene. Consequently, the spatial features encoded by different model branches could be fused to reduce the number of the parameters and improve scalability. This approach, defined as continual dynamic neural graphics primitives (CD-NGP), is designed to handle the dynamic nature of scenes in the online setting. Equipped with the hash encoders and MLP regressors defined in Section 2.2, the $k$-th branch of the model contains spatial hash encoder $\Psi_{k}(\mathbf{x})$, temporal hash encoder $\Gamma_{k}(t)$, and MLPs $f_{\theta_{1k}},f_{\theta_{2k}}$. Please note the dimensions of the features is independent of the sizes of the hash tables: a larger hash table alleviates hash collision at the cost of more storage consumption, but the features encoded by different sizes of hash tables share the same size, i.e. $\Psi_{k}(x)\in\mathbf{R}^{LF}$ holds for all $k$. To alleviate hash collision and reduce model size simultaneously, we apply an asymmetric hash table design: the hash table in the first branch (base branch) is set as $2^{P_{1}}$, and the hash tables in the subsequent branches (auxiliary branches) are set as $2^{P_{2}}$, where $P_{1}$ is much larger than $P_{2}$. The base branch is designed to reconstruct the scenes with initial movements and static objects, and the auxiliary branches are designed to compensate for residuals. The re-use of the features in the base branch ensures high scalability and alleviates the catastrophic forgetting problem. When ($P_{1},P_{2})=(19,14)$, the hash table in the second branch consumes only $1/32$ of the parameters in the first branch. We substitute the $\Psi(\mathbf{x})$ in Eq. 3 with the fused feature to obtain the volume density $\sigma$ at the $k$-th model branch under the continual learning setting, defined as:
$(\sigma,\mathbf{H})=f_{\theta_{1k}}((\Psi_{0}(\mathbf{x})+\Psi_{k}(\mathbf{x}))\oplus\Gamma_{k}(t)),$ (6)
where $\Psi_{0}(\mathbf{x})$ denotes the spatial encoder of the base branch, $\Psi_{k}(\mathbf{x})$ denotes the spatial encoder of the current branch. The view-dependent colors are regressed as in Eq. 4 with the color MLP $f_{\theta_{2k}}$. To cope with large motions and complex transient effects in the long time intervals, we assign $T_{episode}$ chunks as an episode. The MLP regressors are re-trained from initialization in the first branch of each episode to adapt to recent changes. We further illustrate the whole process of our method in Algorithm 1, where $\eta_{init}$ and $\eta_{aux}$ denote the numbers of iterations of the initial branch and the auxiliary branches, respectively. $\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)}$ denote the model parameters in the branches.
Algorithm 1 Continual learning pipeline
0:  $N_{chunk},T_{chunk},T_{episode},\eta_{init},\eta_{aux},\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)},i\in\mathbb{N}$
0:  The whole representation for dynamic scenes.
[…]
2:  while $k$<$N_{chunk}$ do
[…]
5:     if $k\leavevmode\nobreak\ \mathbf{mod}\leavevmode\nobreak\ T_{episode}\leavevmode\nobreak\ =0$  then
[…]
7:     else
[…]
10:     end if
11:     Train the $k$-th model branch $(\Psi_{k},\Gamma_{k},\theta_{1}^{(k)},\theta_{2}^{(k)})$ for $\eta$ steps.
[…]
13:  end while
The scalability of our method derives from the redundancy of dynamic scenes and the flexibility of hash table sizes. As pointed out by [45], most points ($>90\%$ in DyNeRF dataset by their observation) in the dynamic scenes are static. We thus leverage the shared static points across time and represent the first chunk with a large hash table and the subsequent chunks with small hash tables to ensure scalability. Since the spatial hash tables occupy the most parameters in the model, and most chunks are encoded by the subsequent branches, our asymetric hash table design enables continual learning of dynamic scenes and high scalability in streaming scenarios simultaneously. The total bandwidth consumption of our method is $\mathcal{O}(N_{chunk}\cdot 2^{P_{2}}LF+2^{P_{1}}LF)$, while the previous method [43] requires offline training and consumes $\mathcal{O}(\mathbf{K}\cdot 2^{P_{1}}LF)$ bandwidth, where $\mathbf{K}$ ranging from $0.5$ to $16$ is their bandwidth budget controlling the bitrates in streaming scenarios. If the bandwidth is extremely limited, the hash tables in our model could be arranged as a combination of 3D representations and 2D representations like [38] to reduce the bandwidth cost additionally.
2.4 Rendering and optimization
Given the sampled optical attributes on the rays in Eq. 1, we adopt volume rendering to get the 2D pixels, defined as:
$\hat{C}(r)$ $=\sum_{i=1}^{N}T_{i}\alpha_{i}\mathbf{c_{i}}=\sum_{i=1}^{N}T_{i}(1-e^{-\sigma_{i}\delta_{i}})\mathbf{c_{i}},$ (7)
where $T_{i}=\exp(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j})$ denotes the transmittance, $\delta_{i}$ is the interval between neighboring samples, $N$ is the number of samples along the ray, $\mathbf{c_{i}}$ is the output of Eq. 4. The model is optimized by L2 photometric loss with three regularization terms, defined as:
$L=\frac{1}{|\mathbf{R}|}\sum_{r\in\mathbf{R}}\|C(r)-\hat{C}(r)\|_{2}^{2}+\lambda_{d}L_{d}+\lambda_{o}L_{o}+\lambda_{r}L_{r},$ (8)
where $\mathbf{R}$ denotes the set of the camera rays. $L_{d}$ is the distortion loss in Mip-NeRF 360 [3], defined as:
$L_{d}$ $=\sum_{i,j}w_{i}w_{j}\left\lvert\frac{s_{i}+s_{i+1}}{2}-\frac{s_{j}+s_{j+1}}{2}\right\rvert$
$+\frac{1}{3}\sum_{i}w_{i}^{2}(s_{i+1}-s_{i}),$ (9)
where $s_{i}$ denotes the normalized ray distance along the ray, $w_{i}=T_{i}\alpha_{i}$ is the weight of the $i$-th sample.
$L_{o}$ in Eq. 8 is the entropy on the opacity [25], defined as:
$L_{o}=-o\log(o),\quad o=\sum_{j=1}^{N}T_{j}\alpha_{j},$ (10)
$L_{r}$ in Eq. 8 is the L1 regularization on the spatial features in the current branch as:
$L_{r}=||\Psi_{k}(\mathbf{x})||_{1}.$ (11)
We set $\lambda_{r}$ as 0.001 to encourage using the features encoded by the base branch. Based on [25], we set a large value of $\lambda_{d},\lambda_{o}$ as 0.005 to alleviate foggy effects.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Example frames sampled from the proposed long multi-view video dataset.
3 Experiments
3.1 Datasets
The DyNeRF dataset [28] is the most prevalent benchmark for dynamic view synthesis, providing high-quality multi-camera videos on which current state-of-the-art methods are primarily evaluated. This dataset includes six unique cooking scenes, each featuring varying illumination, food topological changes, and transient effects such as flames. It consists of five videos, each containing 300 frames, and one extended video with 1200 frames, captured at a resolution of $2704\times 2028$ with up to 21 camera views at 30 frames per second (FPS). Following the approach of HexPlane [7], we excluded the unsynchronized scene coffee martini and conducted experiments on the remaining scenes, using the first 300 frames of the 1200-frame video flame salmon for evaluation.
Meetroom dataset [27] is another dataset for dynamic view synthesis that contains 3 scenes captured by 13 synchronized cameras at $1280\times 720$ resolution. The scenes are 300 frames long and captured at 30 FPS.
To further validate the effectiveness of our method on long video sequences, we introduce a new dataset comprising 5 dynamic scenes at a resolution of $3840\times 2160$ as shown in Fig. 2. Similar to those in the DyNeRF dataset [28], these scenes are captured by 20 GoPro cameras operating at 30 FPS. The cameras are arranged in two rows and placed in an arc facing the scenes. The camera in the center of the upper row is designated for validation. We synchronize the multi-view videos and incorporate chessboard images within the scenes to facilitate camera calibration and pose estimation. Our dataset features more challenging dynamics, including significant non-rigid human motions and transient effects like screen slides, compared to the DyNeRF dataset. To comprehensively assess the scalability of our method, we provide 5 scenes with a duration exceeding 200 seconds (6000 frames), which is over 10 times longer than the previous DyNeRF dataset. Our dataset is collected to evaluate the scalability of our method and to support further research in the community. We have carefully reviewed the videos to ensure they do not contain any controversial or offensive content before using them for research purposes. While the faces in the videos may contain identifiable information, this information will be used solely for research and will not be disclosed to any third party. All videos were collected with the explicit consent of all participants.
3.2 Implementation details
Our implementation is based on HexPlane [7], Instant-NGP [31], and NGP-pl [25]. All the metrics are measured on a single RTX 3090 GPU if not otherwise specified. Similar to the previous work [7, 28], we select the view in the center with index 00 for evaluation on both datasets. Following most of the state-of-the-art methods [28, 27, 17, 45, 43], we train and evaluate the model at resolution $1352\times 1014$ on the DyNeRF dataset. We follow [27] and use $1280\times 720$ resolution on the Meetroom dataset. We use $1280\times 720$ resolution for our dataset to fit into the host memory. We use the importance sampling approach described in [28] at a batch size of 1024 rays. The model parameters are optimized by Adam [23], with hyper parameters $\epsilon=10^{-15},\beta=(0.9,0.96)$ for stability. The learning rate is set as 0.001 and scheduled by cosine function. We set the hash table sizes as $2^{P_{1}}=2^{19}$ in the first branch and $2^{P_{2}}=2^{14}$ in the subsequent branches. The structural parameters $(L,F)$ in the spatial hash encoders are set as $(12,2)$ if not otherwise specified. At the base branch, we initialize the models and train for $\eta_{init}=18k$ iterations, and for each of the subsequent branches we train for $\eta_{aux}=3k$ iterations using the previous branch for initialization for the DyNeRF dataset and our dataset. We use $\eta_{init}=33k$ and $\eta_{aux}=3k$ for the Meetroom dataset. We use LeakyReLU activation in the MLP regressors. Strategies for updating the occupancy grids are identical with [31]. To save storage, we use only one shared occupancy grid for all the branches. We report the model sizes with the occupancy grid under bfloat16 precision and parameters in float32 precision. The episode length $T_{episode}$ is set as $30$ only for videos longer than 300 frames in our own dataset. We set the dimensions of the latent feature $\mathbf{H}$ in Eq. 3 as 48. We use MLPs with two hidden layers and the dimensions of the hidden layers in Eqs. 3 and 4 are set as 128, and 64 respectively. We use MLPs with identical structures for different branches. We set the resolutions of the hash encoders of the CD-NGP within $(16,2048)$ times the scale of the scene.
3.3 Results and scalability analysis
Table 1: Comparisons between our CD-NGP and state-of-the-art methods on the DyNeRF dataset. Memory denotes the maximum host memory required during training. $\dagger$ denotes the methods evaluated only on the flame salmon scene. - denotes the metrics not reported in the original publications.
[…]
DyNeRF [28] $\checkmark$ 300 29.58 0.020 0.083 7 days 28MB >100GB
NeRF-T [28] $\checkmark$ 300 28.45 0.023 0.10 - - >100GB
Interp-NN [34] $\checkmark$ 300 29.88 - 0.096 2 days 20MB >100GB
HexPlane [7] $\checkmark$ 300 31.57 0.016 0.089 96 min 200MB 62GB
K-Planes [17] $\checkmark$ 300 31.63 - - 110 min 200MB >100GB
MixVoxels [45] $\checkmark$ 300 30.71 0.024 0.162 15 min 500MB >100GB
4DGS [55] $\checkmark$ 300 32.01 - 0.055 6 hours 4000MB <10GB
HyperReel [1] $\checkmark$ 50 31.1 - 0.096 9 hours 360MB 18GB
NeRFplayer [43] $\checkmark$ 300 30.29 - 0.152 5.5 hours - >100GB
StreamRF [27] $\checkmark$ 1 28.85 0.042 0.253 75 min 5000MB <10GB
HexPlane-PR $\checkmark$ 10 24.03 0.081 0.244 100 min 318MB 14GB
HexPlane-GR $\checkmark$ 10 25.17 0.050 0.272 125 min 318MB 14GB
INV [48] $\checkmark$ 1 29.64 0.023 0.078 40 hours 336MB <10GB
CD-NGP (ours) $\checkmark$ 10 30.23 0.027 0.198 75 min 113MB 14GB
3.3.1 Comparisons on DyNeRF dataset.
Table 1 shows the quantitative comparisons between CD-NGP and other state-of-the-art methods on DyNeRF dataset. The methods in [28, 34] are trained on 8 V100 GPUs or equivalently 4 A100 GPUs. The methods in [28, 34, 43] have no official implementations available. We are unable to obtain the missing quality metrics in [17] because of their huge memory occupation in the default resolution. We report the metrics in the publication if not otherwise specified. HexPlane [7] uses resolution $1024\times 768$ in their implementation, so the memory occupation is much smaller than other offline methods. We use the official implementation of StreamRF [27] and INV [48] and report the average metrics on the 5 synchronized scenes of the dataset. StreamRF [27] and INV [48] are trained per frame. NeRFplayer [43] is trained offline but streamable upon inference. 4DGS [55] caches the frames onto disk for training to alleviate memory occupation with a trade-off on training speed. The reconstruction quality is measured by PSNR for pixel-wise similarities, DSSIM for structural similarities, and LPIPS [57] for perceptual similarities (by AlexNet [24]) following common practices [28, 7, 45]. To prove the viability of our parameter-isolation-based method, we apply the previous state-of-the-art method HexPlane [7] to continual learning setting with parameter-regularization (HexPlane-PR) and generative replay (HexPlane-GR) as additional baselines.
We compare our method with both online and offline baselines. As demonstrated in Table 1, our CD-NGP is 32×32$\times$32 × faster than the MLP-based continual representation INV [48], $\times4$ smaller than the voxel-based continual representation StreamRF [27] and significantly outperforms the two additional baselines in the continual learning setting. The undesirable results from HexPlane-PR and HexPlane-GR show the limitations of the regularization-based and replay-based methods in dynamic scenes: unlike the static scenes, the objects are not stationary across time and the models with limited complexity are not capable of reconstructing unlimited transient effects. Benefiting from the parameter isolation like StreamRF [27] and INV [48], our method achieves fast convergence and compact model size simultaneously. It leverages the hash tables to ensure fast convergence and reuses the features from differently configured spatial hash tables to ensure high scalability. Compared with prevailing offline methods [28, 34, 7, 17, 45, 55], our method only requires caching $T_{chunk}=10$ frames into the buffer by continual learning and thus reduces memory occupation by $85\%$. Our method also reduces $40\%$ less model size (to $0.4$MB/frame) by leveraging the feature fusion of the base and the auxiliary branches and delivers comparable quality to the offline methods.
(a) HexPlane-offline
(b) MixVoxels
(c) StreamRF
(d) HexPlane-PR
(e) HexPlane-GR
(f) CD-NGP (ours)
Figure 3: Comparison of reconstruction quality on the most challenging flame salmon scene in DyNeRF dataset. PSNRs are computed under the resolution $1352\times 1014$.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: The results of the proposed CD-NGP on DyNeRF dataset.
Figure 4 provides qualitative comparisons. Compared with offline methods in Fig. 3(a), Fig. 3(b) and the online method in Fig. 3(c), our model is much smaller without large quality loss. Compared with regularization-based and replay-based online methods in Fig. 3(d) and Fig. 3(e) respectively, our method achieves significantly higher quality. Results of other scenes including depth maps are shown in Fig. 4.
(a) GT
(b) HexPlane
(c) MixVoxels
(d) INV
(e) StreamRF
(f) Ours
Figure 5: Comparison of reconstruction quality on the trimming scene in the Meetroom dataset. Our method produces comparable quality to both offline and online state-of-the-art methods but with much less memory and storage cost.
3.3.2 Comparisons on the Meetroom dataset.
Table 2 and Fig. 5 show the quantitative and qualitative comparisons on the meetroom dataset. Fig. 6(a) is the PSNR curve of the compared methods on the trimming scene. We use the released configuration of StreamRF [27] for the meetroom dataset to obtain the results. As shown in Table 2, HexPlane-GR and HexPlane-PR produce significantly lower quality than the other methods on the meetroom dataset, we thus exclude them from Fig. 6(a) for better visualization. The metrics in Table 2 and the curves in Fig. 6(a) show the advantage of our method over both online and offline baselines. It outperforms the implicit MLP-based online method [48] regarding reconstruction quality, speed, and model size. Moreover, our method consumes only a fraction of storage than the explicit voxel-based online baseline [27] while maintaining comparable reconstruction quality as compared in Fig. 5. When compared with offline baselines, our method produces comparable quality to MixVoxels [45] and outperforms HexPlane [7] in terms of quality and speed by much less memory cost.
Figure 6: Combined PSNR curves for various scenes and methods. (a) The trimming scene in the Meetroom dataset. Solid and dotted lines denote offline and online methods respectively. (b) The first 1200 frames of seminar scene in our dataset. The solid lines and dotted lines denote offline and online methods respectively. (c) The first 2400 frames of seminar scene in our long video dataset. (d) the sear steak scene in the DyNeRF dataset. Solid and dotted lines denote symmetric and asymmetric setting, respectively.
Table 2: Comparisons of the metrics on the Meetroom dataset.
[…]
Mixvoxels $\checkmark$ 26.96 0.027 0.103 15min 500MB
HexPlane $\checkmark$ 24.71 0.043 0.221 100min 200MB
HexPlane-PR $\checkmark$ 18.63 0.112 0.359 100min 318MB
HexPlane-GR $\checkmark$ 19.49 0.099 0.400 100min 318MB
StreamRF $\checkmark$ 26.48 0.029 0.170 75min 1710MB
INV $\checkmark$ 23.87 0.041 0.120 40hours 336MB
CD-NGP (ours) $\checkmark$ 26.01 0.039 0.191 70min 113MB
3.3.3 Spatial representations and scalability analysis.
Our proposed continual learning framework is robust against changes in spatial backbones, including spatial hash encoder configurations and feature fusion methods. The hash encoder in Eq. 5 could be changed into the hybrid representation (MERF) [38] and the plane representation like those in [17, 21]. The MERF representation is defined as:
$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}\leavevmode\nobreak\ P_{m}(x,y,z)+Q_{mxy}(x,y)+Q_{myz}(y,z)+Q_{mzx}(z,x),$ (12)
where $Q_{mij}(i,j)$ denotes the bi-linearly interpolated features from the 4 nearby vertices in the orthogonally projected planes w.r.t. axes $(i,j)$ at the $m$-th resolution. The plane representation is defined as:
$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}Q_{mxy}(x,y)+Q_{myz}(y,z)+Q_{mzx}(z,x).$ (13)
For the simplicity of the notations, we denote the models adopting Eq. 5, Eq. 12, and Eq. 13 as CD-NGP, CD-MERF, and CD-Plane respectively.
We compare the results of different spatial representations on the DyNeRF dataset and list the parameter numbers in Table 3. The alternatives of spatial representations include CD-MERF and CD-Plane. The element-wise product for plane representation in [17] produces unsatisfactory results in the experiments, we thus remove the $P_{m}(x,y,z)$ in Eq. 12 and fuse the features $Q_{mij}(i,j)$ by concatenation in the implementation. As shown in Table 3, the hybrid representation CD-MERF achieves high compression and comparable model sizes to methods for static scenarios [31, 21] ($\sim$60MB) without significant quality loss. We also adopt the 3D+4D representation similar to that in [34] for comparison. This representation is implemented by concatenating the spatial $(x,y,z)$ features and spatial-temporal $(x,y,z,t)$ features encoded by a 4D extension of Eq. 5, i.e. $\Gamma(x,t)=\bigoplus_{m=1}^{L}P_{m}(x,y,z,t)$. However, the asymmetric hash sizes (using $(P_{1},P_{2})=(19,14)$) in the 3D+4D representation results in serious feature collisions and deliver bad results (PSNR < 16 dB on 3 scenes). We thus set $(P_{1},P_{2})=(14,14)$ for the 3D+4D representation in Table 3 to keep the size of the model close to our CD-NGP for fair comparison. The metrics show the advantage of our separated encoding of spatial-temporal $(x,y,z,t)$ features explained in Eqs. 3 and 6: representing the features of spatial coordinates $(x,y,z)$ and temporal coordinate $(t)$ separately avoids feature collisions and lead to better results.
The results in Table 3 also show the flexibility in the feature fusion methods. We replace the element-wise sum in Eqs. 6 and 12 with vector concatenation and denote them as CD-NGP-C and CD-MERF-C in Table 3. Benefiting from larger feature dimensions, these methods lead to slightly better quality (especially when measured in LPIPS) at the cost of more training time, and the reasonable trade-off also shows the robustness of our method.
Table 3: Quality metrics, parameter numbers, and bandwidth costs of different spatial representations. The parameter numbers of each module of the 30-branch representations are also shown. #$\#3DEnc.$ and #$\#2DEnc.$ denote the parameter numbers of the 3D encoders and 2D encoders in the base branch and one auxiliary branch accordingly. $\mathbf{B_{min}}$ and $\mathbf{B_{avg}}$ are minimum and average bandwidth costs in MB/frame. ##\## denotes the number of parameters per branch (10 frames).
[…]
CD-NGP 21M 30.23 0.198 75min 113MB $(9.9M,0.37M)$ $0.39M$ 0.16 0.38
3D+4D 23M 29.31 0.242 82min 122MB $(0.73M,0.73M)$ $0.76M$ 0.31 0.41
CD-NGP-C 21M 30.52 0.174 80min 114MB $(9.9M,0.37M)$ $0.40M$ 0.16 0.38
CD-Plane 14M 29.16 0.215 77min 83MB - $(1.9M,0.082M)$ $0.28M$ 0.11 0.28
CD-MERF 7.5M 29.93 0.201 109min 57MB $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.16M$ 0.061 0.19
CD-MERF-C 7.5M 30.22 0.174 115min 57MB $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.17M$ 0.065 0.19
The parameter numbers and bandwidth consumption in Table 3 also prove the superior scalability of our method, where 3D Enc. and 2D Enc. denote the spatial encodings according to voxel and plane indices respectively. Following MERF [38] for static scenes, we use a smaller hash table for the hybrid representation in Eq. 12 than pure voxel hash grid [31] representation in Table 3 to ensure the compactness and scalability of our method. Denoting the lengths of hash tables along pure voxels in a branch of CD-NGP as $2^{P}$, the lengths of hash tables in the CD-MERF are set as $2^{P-3}$ and $2^{P-4}$ for voxel and plane encoders respectively. The number of the parameters of the MERF representation is reduced to $1-2^{-3}-3\times 2^{-4}=31.25\%$ of the voxel hash grids. Since the hash tables comprise most of the parameters, the model size is reduced from $21M$ to $7.5M$ as listed in Table 3. Likewise, we use $2^{P-2}$-length hash tables for the Plane representation in Table 3 to balance the quality and the model size. The structural configurations $(L,F)$ of the plane representation in Table 3 are set as $(6,4)$ , and the resolution of the hash encoders of the CD-Plane are set within $(64,2048)$ for better results. Following [27, 43, 1, 48] we also report the minimum per-frame bandwidth cost $\mathbf{B_{min}}$, which only depends on the parameter numbers of the auxiliary branches and is thus significantly lower than the average bandwidth consumption $\mathbf{B_{avg}}$. Since the base encoder and the density grid could be shared among different branches and transmitted only once in the beginning, the minimum bandwidth consumption is reduced to $0.065\sim 0.16$ MB per frame, which is 100×100$\times$100 × less than the voxel-based StreamRF [27] and 7×7$\times$7 × less than the pure MLP representation INV [48].
3.3.4 Long sequence dataset and scalability analysis.
As introduced in Section 2.3, our method shows superior scalability: the total size of the model is $\mathcal{O}(N_{chunk}\cdot 2^{P_{2}}LF+2^{P_{1}}LF)$, where $N_{chunk}$ is the number of chunks in the sequence. To prove the scalability of our method, we provide quantitative comparisons on our proposed long multi-view video dataset in Table 4 and Fig. 6(b) to further show the superior scalability of our method. We split the sequence into 300-frame chunks for the compared methods since the offline baseline [7] is designed for 300-frame sequences in the DyNeRF dataset. We set the hash table size as $2^{12}$ in the auxiliary branches and use $T_{chunk}=5$ in our model. As shown in Table 4, our method reduces the memory and storage consumption by at least 4×4$\times$4 × with comparable quality to HexPlane [7]. Compared with the results in Table 1, the performance gap in storage consumption is even larger in long video sequences. This is because the majority of the parameters in our method are used for the auxiliary branches, and the base branch only takes up a small fraction of the parameters. Consequently, applying smaller hash tables in the auxiliary branches and leverage the re-use of the features in the base branch leads to a significant reduction in the model size.
The PSNR curves in Fig. 6(b) further demonstrate the stability of our method. The compared regularization-based and replay-based methods suffer from catastrophic forgetting: the PSNRs at the beginning of the chunks are much lower than those in the end. The problem is properly eliminated by our method through parameter isolation, and our method continuously produces high-quality results for the whole sequence. We also provide comparisons between the rendered results in Fig. 7. The regularization method displayed in Fig. 7(a) reconstructs the TV screen in the background erroneously. The replay method displayed in Fig. 7(b) fails to recover the details. Ours in Fig. 7(f) produces comparable quality to the offline baseline in Fig. 7(c).
To further demonstrate the scalability of our method, we conduct experiments on the first 2400 frames of the seminar scene in our long video dataset. We further show the PSNR curves in Fig. 6(c) and compare the model sizes in Fig. 8. Our method produces high-quality results while consuming only a fraction of storage space compared with the other online baselines.
3.4 Ablation studies
(a) HexPlane-PR
(b) HexPlane-GR
(c) HexPlane
(d) INV
(e) StreamRF
(f) CD-NGP (ours)
Figure 7: Comparison of reconstruction quality on the 19th frame of the seminar scene in our long video dataset.
Figure 8: Comparison of the model sizes of our method and other online baselines. Our method shows superior scalability over others.
3.4.1 Lengths of hash tables.
Benefiting from the multi-branch hash table design, our model is highly efficient and configurable. We change the lengths of the hash tables of the initial branch and the auxiliary branches and obtain the results on the DyNeRF dataset in Table 5. $(P_{1},P_{2})$ are the log2 values of the hash table lengths of the base branch and the auxiliary branches. The whole model contains hash encoders, MLP regressors, and the occupancy gridd, the sizes of the models are thus larger than the value deduced from the combinations of the hash table and MLP regressors. Although larger hash tables lead to less hash collisions, the default configuration $(19,14)$ in CD-NGP produces better results than the naive $(16,16)$ configuration even if the storage consumption is $40\%$ lower. Moreover, the $(19,12)$ configuration of CD-NGP significantly outperforms the naive $(14,14)$ configuration at a similar storage cost. To further demonstrate the superiority of the asymmetric hash table design and prove the scalability of our method, we provide the PSNR curves of different hash table designs on the sear steak scene in Fig. 6(d). Equipped with a larger hash table in the base branch, the asymmetric setting $(19,14)$ and $(19,12)$ outperforms others in the beginning and produces comparable metrics to the naive $(16,16)$, $(14,14)$ configuration respectively in the subsequent chunks. The $(18,18)$ in Table 5 configuration delivers the best results (30.55 dB in PSNR) on the DyNeRF dataset, but at the cost of an exceptionally large model size (638MB). However, the CD-NGP-C configuration with $(P_{1},P_{2})=(19,14)$ listed in Table 3 achieves a comparable PSNR (30.52 dB) with a much smaller model size (114MB, $80\%$ compression). The advantage comes from our re-use of the spatial features encoded by different branches: the base branch with a large hash table recovers more environment details, and the auxiliary branches could thus use fewer parameters to cope with recent changes only. The combination of the base encode and the auxiliary encoders lead to a more efficient and scalable model.
Table 4: Quantitative comparisons between our CD-NGP and baselines on the long video sequences in our dataset. Metrics are measured on the first 1200 frames and averaged among the different scenes. The compared HexPlane baselines are trained with a chunk size of 300 frames.
[…]
MixVoxels $\checkmark$ 26.94 0.022 0.118 2000MB 80 min 80GB
HexPlane $\checkmark$ 27.74 0.022 0.117 800MB 8 hours 74GB
HexPlane-PR $\checkmark$ 20.90 0.095 0.286 1272MB 8 hours 12GB
HexPlane-GR $\checkmark$ 24.82 0.052 0.282 1272MB 8 hours 12GB
StreamRF $\checkmark$ 24.72 0.036 0.273 13GB 4 hours 12GB
INV $\checkmark$ 27.12 0.025 0.086 1344MB 160 hours 12GB
CD-NGP (ours) $\checkmark$ 27.15 0.039 0.188 201MB 10 hours 12GB
Table 5: Ablation study on hash table sizes. denotes the default setting.
[…]
$(18,18)$ 151M 30.55 0.024 0.161 638MB
$(16,16)$ 41M 30.08 0.027 0.194 191MB
$(14,14)$ 12M 29.32 0.035 0.248 75MB
$(19,14)^{{\dagger}}$ 21M 30.23 0.027 0.198 113MB
$(16,14)$ 13M 29.78 0.030 0.228 79MB
$(19,12)$ 14M 30.09 0.028 0.209 82MB
3.4.2 Temporal representations.
Our continual learning pipeline is also compatible with a variety of representations for the time axis. Currently, there are 2 other prevailing representations for the time axis alone: pure frequency encoding [36] and MLP with frequency-encoding [16]. We replace the temporal hash encoding in our method with the ones mentioned above on DyNeRF dataset and show the results in Table 6. Our time latent code slightly outperforms others on DyNeRF dataset. The robustness against changes in the temporal representations also shows the viability of the continual learning method.
Table 6: Ablation study on temporal representations.
[…]
CD-NGP 30.23 0.0269 0.198
Time-MLP 30.20 0.0273 0.202
Time-Freq 30.19 0.0272 0.203
Table 7: Ablation study on the field composition methods.
[…]
CD-NGP 30.23 0.0269 0.198
Full composition 29.24 0.0367 0.228
Static-dynamic 28.27 0.0568 0.291
3.4.3 Field composition methods.
We also conduct ablation studies on the field composition methods and show the results in Tables 7 and 9. The field composition process is defined as:
$(\sigma,\mathbf{c})=(\sigma_{i}+\sigma_{a},\mathbf{c}_{i}+\mathbf{c}_{a}).$ (14)
Full composition in Table 7 denotes the naive composition of 2 fields from the base branch and the current branch respectively, i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ are the field from the base branch and $\sigma_{a}$ and $\mathbf{c}_{a}$ in Eq. 14 are the field from the current branch. Both branches are encoded with the same architecture as Eq. 3. Static-dynamic in Table 7 denotes a naive composition of the field from a static branch by the base encoder and a dynamic branch by the auxiliary encoders similar to [45], i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ in Eq. 14 are encoded without the $\Gamma(t)$ in Eq. 3. The online training strategies remain the same. Our CD-NGP outperforms the other methods in all metrics. Furthermore, as compared in Fig. 9, the compared composition methods in Figs. 9(c) and 9(d) suffer from salt and pepper noise because of the explicit composition, while our CD-NGP circumvents the problem by fusing features in the latent space and produces better results.
(a) Ground truth
(b) CD-NGP (ours)
(c) Full composition
(d) Static-dynamic
Figure 9: Comparison of different field composition methods.
Table 8: Ablation study on the chunk sizes. The iteration numbers of each chunk are held constant under different $T_{chunk}$ settings.
[…]
30 30.31 0.02654 0.1842 80MB
15 30.54 0.02557 0.1816 97MB
10 30.76 0.02440 0.1793 113MB
5 30.99 0.02268 0.1674 162MB
Table 9: Ablation study on the chunk sizes. The total iterations of the whole 300-frame video are held constant under different $T_{chunk}$ settings.
[…]
30 30.84 0.02392 0.1505 80MB
15 30.76 0.02413 0.1684 97MB
10 30.76 0.02440 0.1793 113MB
5 30.39 0.02602 0.2075 162MB
3.4.4 Chunk size.
We also conduct ablation studies on the chunk sizes and show the results in Tables 8 and 9. We only report the results on the cook spinach scene in the DyNeRF dataset for brevity. As in Table 8, the smaller chunk size leads to better results when controlling the iteration numbers of each chunk. When the total iterations of the whole 300-frame sequence are held constant, the larger chunk size leads to better results as in Table 9. The selection of the chunk size is a trade-off between the quality, robustness, training time, and host memory occupation and should depend on the data, and our method is robust within a wide range of chunk sizes. We select $T_{chunk}=10$ as the default configuration.
4 Conclusion and Limitation
We propose CD-NGP, a continual learning framework for multi-view dynamic scenes based on parameter isolation, along with a self-collected dataset for novel view synthesis under multi-view long-view scenarios. Our proposed method splits the input multi-view videos into different chunks, trains different model branches, and performs reconstruction accordingly. Each branch encodes spatial and temporal features separately using the multi-resolution hash encoding structure. It leverages the redundancy of the dynamic scenes and reuses the spatial features in the first branch to achieve fast convergence, high fidelity, and superior scalability. On the prevailing DyNeRF dataset, it consumes $40\%$ less bandwidth and $85\%$ less memory than recent explicit offline baselines, achieves comparable quality, and maintains the speed advantage over pure-MLP-based offline methods. Compared with online baselines, it delivers state-of-the-art quality and achieves a new balance among reconstruction quality, training speed, bandwidth occupation, and memory cost. On the long video dataset, our method delivers high quality and achieves fast convergence with only a fraction of the storage cost. However, like most NeRF-based methods, it still requires querying the field function and is not suitable for real-time rendering. Further efforts could be made to apply baking methods or Gaussian point clouds to continual learning, enabling real-time rendering in long video sequences for wide applications.
Appendix A Per-scene results
We list the per-scene results of the proposed continual learning method with different spatial representations on the DyNeRF dataset in Table 10, and the per-scene results of the methods in Table 4 on our self-collected long video dataset in Table 12.
Table 10: Detailed results of the proposed CD-NGP and other hash-based methods on the DyNeRF dataset.
Model Flame Salmon Cook Spinach Cut Roasted Beef
[…]
CD-NGP 26.37 0.046 0.289 30.76 0.024 0.179 31.03 0.024 0.180
3D+4D 25.72 0.059 0.355 29.73 0.032 0.220 30.06 0.031 0.224
CD-NGP-C 26.65 0.042 0.268 31.03 0.022 0.159 31.26 0.022 0.158
CD-Plane 25.48 0.060 0.317 29.41 0.034 0.201 30.11 0.033 0.195
CD-MERF 26.15 0.049 0.292 30.35 0.029 0.186 30.71 0.028 0.186
CD-MERF-C 26.56 0.044 0.255 30.69 0.026 0.158 31.04 0.025 0.163
Model Flame Steak Sear Steak Average
[…]
CD-NGP 31.27 0.021 0.174 31.71 0.019 0.170 30.23 0.027 0.198
3D+4D 30.33 0.028 0.205 30.73 0.026 0.207 29.31 0.035 0.242
CD-NGP-C 31.64 0.018 0.140 32.02 0.017 0.142 30.52 0.025 0.174
CD-Plane 30.26 0.029 0.180 30.56 0.027 0.181 29.16 0.037 0.215
CD-MERF 31.01 0.025 0.172 31.44 0.022 0.170 29.93 0.031 0.201
CD-MERF-C 31.13 0.024 0.151 31.66 0.020 0.141 30.22 0.028 0.174
Table 11: Per-scene results of the compared methods on the meetroom dataset in Table 2.
Model Trimming Discussion
[…]
MixVoxels 27.21 0.026 0.093 27.42 0.026 0.110
HexPlane 22.13 0.062 0.263 26.28 0.032 0.212
HexPlane-PR 17.30 0.140 0.404 18.53 0.111 0.355
HexPlane-GR 12.80 0.184 0.554 23.31 0.061 0.330
StreamRF 26.72 0.029 0.170 26.65 0.028 0.169
INV 27.08 0.026 0.089 24.57 0.036 0.110
CD-NGP 27.58 0.040 0.164 26.01 0.040 0.214
Model VRheadset Average
[…]
MixVoxels 26.25 0.029 0.107 26.96 0.027 0.103
HexPlane 25.72 0.035 0.187 24.71 0.043 0.221
HexPlane-PR 20.05 0.085 0.317 18.63 0.112 0.359
HexPlane-GR 22.37 0.054 0.316 19.49 0.099 0.400
StreamRF 26.06 0.029 0.171 26.48 0.029 0.170
INV 19.95 0.061 0.160 23.87 0.041 0.120
CD-NGP 24.43 0.037 0.193 26.01 0.039 0.191
Table 12: Detailed results of the proposed CD-NGP and other hash-based methods on our long video dataset.
Model Demo Discussion Kungfu
[…]
MixVoxels 27.50 0.021 0.103 28.01 0.016 0.097 24.43 0.033 0.140
HexPlane 28.92 0.016 0.095 27.17 0.024 0.136 25.14 0.032 0.136
HexPlane-PR 21.10 0.090 0.268 20.98 0.094 0.315 18.85 0.112 0.309
HexPlane-GR 24.85 0.041 0.243 24.03 0.050 0.296 22.15 0.072 0.336
StreamRF 24.86 0.032 0.233 24.41 0.029 0.258 23.75 0.043 0.301
INV 26.55 0.025 0.093 28.51 0.017 0.054 24.57 0.037 0.120
CD-NGP 26.66 0.030 0.161 29.66 0.057 0.137 23.66 0.055 0.257
Model Reading Seminar Average
[…]
MixVoxels 25.99 0.021 0.144 28.78 0.018 0.108 26.94 0.022 0.118
HexPlane 27.88 0.018 0.099 29.59 0.018 0.120 27.74 0.022 0.117
HexPlane-PR 21.85 0.087 0.287 21.78 0.092 0.280 20.91 0.095 0.292
HexPlane-GR 26.52 0.040 0.266 25.77 0.053 0.306 24.66 0.051 0.289
StreamRF 24.06 0.040 0.325 26.53 0.037 0.250 24.72 0.036 0.273
INV 26.39 0.026 0.106 29.60 0.018 0.054 27.12 0.025 0.085
CD-NGP 26.70 0.031 0.212 28.97 0.024 0.169 27.13 0.039 0.201